OpenStudy (mooredog):

The student council wants to rent a ballroom for the junior prom. The ballroom's rental rate is $1500 for 3 h and $125 for each additional half hour. Suppose the student council raises $2125. What is the maximum number of hours for which they can rent the ballroom. I just the need the inequality for this word problem, I can solve it by myself. I just don't know how to get the inequality.

OpenStudy (jackthegreatest):



OpenStudy (jackthegreatest):

whoops didnt see the posts bottom half

OpenStudy (jackthegreatest):

dont look at the answer haha

OpenStudy (mooredog):

haha its all good :) thank you so much!

OpenStudy (jackthegreatest):

btw make sure you add h to 3

OpenStudy (mooredog):

huh? do what?

OpenStudy (jackthegreatest):

5 + 3 = 8

OpenStudy (danjs):

It says $125 for each extra Half-Hour , and you can spend up all and equal the 2125. If you use h as the number of extra full hours \[$2125 \ge $1500 + $125*(2*h)\]

OpenStudy (danjs):

\[h \le \frac{ 5 }{ 2 }~hrs\] at most 2.5 extra hours

OpenStudy (jackthegreatest):

oh shoot @DanJS is right

OpenStudy (jackthegreatest):

I did not catch it haha